OpenStudy (he66666):

Trig Identity question? The angle x lies in the interval pi/2≤x≤pi, and sin²x=4/9. Determine cos(x/2). answer: √[(3-√5)/6] How do you solve this question (using trig identities)? It gets confusing because it's x/2, not just x. I keep getting 7/9 as my answer..

OpenStudy (anonymous):

\[cos(\frac{x}{2})=\pm \sqrt{\frac{1+cos(x)}{2}}\]

OpenStudy (anonymous):

square root should be over whole thing.

OpenStudy (he66666):

How does it have a square root? :S Is it derived from a trig identity?

OpenStudy (anonymous):

\[sin^2(x)=\frac{4}{9}\] \[sin(x)=\pm \frac{3}{4}\]

OpenStudy (anonymous):

derived from "double angle" formula \[cos(2x)=2cos^2(x)-1\] replace 2x by x t so x gets replaced by \[\frac{x}{2}\] and then solve for \[cos(\frac{x}{2})\]

OpenStudy (anonymous):

\[cos(x)=2cos^2(\frac{x}{2})-1\] \[cos(x)+1=2cos^2(\frac{x}{2})\] \[\frac{cos(x)-1}{2}=cos^2(\frac{x}{2})\]

OpenStudy (anonymous):

typo should be \[\frac{cos(x)+1}{2}=cos^2(\frac{x}{2})\]

OpenStudy (anonymous):

then take the square root of the whole thing. plus or minus of course. i cannot seem to type set it.

OpenStudy (anonymous):

in case \[sin(x)=\frac{3}{4}\] do you know how to find \[cos(x)\]?

OpenStudy (he66666):

yes, pythagorean theorem?

OpenStudy (anonymous):

whoa hold the phone i am tired. it is \[sin(x)=\frac{4}{9}\]

OpenStudy (anonymous):

damn \[sin^2(x)=\frac{4}{9}\] \[sin(x)=\frac{2}{3}\]

OpenStudy (anonymous):

draw a triangle opposite side 2 and hypotenuse 3 my mistake sorry

OpenStudy (anonymous):

adjacent side is \[\sqrt{3^2-2^2}=\sqrt{9-4}=\sqrt{5}\]

OpenStudy (anonymous):

so \[cos(x)=\frac{\sqrt{5}}{3}\]

OpenStudy (anonymous):

now plug in \[\frac{\sqrt{5}}{3}\] in the formula up top to get your answer. i will write it if you like.

OpenStudy (he66666):

no it's alright, I got it. Thanks so much for the help satellite! :)

OpenStudy (anonymous):

oh another mistake i forgot x was in quad II so cosine is negative. it is \[-\frac{\sqrt{5}}{3}\]

OpenStudy (anonymous):

make sure you use \[-\frac{\sqrt{5}}{3}\]

OpenStudy (he66666):

alright, I will :) thanks

OpenStudy (anonymous):

welcome